# Alpha Berkeley: A Scalable Framework for the Orchestration of Agentic Systems

### v1

@@ IV.1 Wind Farm Example
Consider following user request:
@@
This example shows how our framework automatically decompose this query into a coordinated execution plan that transforms this user request into an transparent, interpretable workflow of steps that are using autonomous agents for temporal reasoning, operating with diverse data sources, use domain expertise and even perform an data analysis, all while maintaining transparency and human oversight necessary for reliable industrial applications.
@@
It demonstrates the framework’s ability to coordinate heterogeneous data sources within a unified workflow. In the workflow, the Steps 2-4 retrieve turbine sensor data, weather measurements, and extract performance benchmarks from technical documentation. The knowledge retrieval step (Step 4) highlights the framework’s ability to integrate external data sources, incorporating databases and knowledge repositories via the data source provider pattern. The knowledge provider retrieves numerical parameters and structured data that are formatted for direct consumption by downstream Python code generation, showcasing integration of external knowledge systems into dynamic analysis workflows.
In order to reliably generate Python code, we do automated hierarchical analysis planning for a given user’s query followed by dynamic code generation. First, language model creates a structured computational plan that breaks the analysis down into discrete phases (data preparation, performance metrics calculation, industry benchmark comparison). This plan is then converted into executable Python code that properly accesses the structured context data from the previous steps.
@@
IV.2 Advanced Light Source Deployment
The ALS at Lawrence Berkeley National Laboratory is a U.S. Department of Energy synchrotron facility that has provided bright X-ray and ultraviolet light to a broad scientific community for more than three decades [42]. In this case study, the Alpha Berkeley Framework is applied in the ALS control room, where the accelerator is operated and optimized to deliver beams to over 40 experimental stations. This environment poses distinctive challenges: real-time integration with control systems, coordination across many heterogeneous subsystems, and strict adherence to safety protocols.
@@ IV.2 Advanced Light Source Deployment
Figure 3 shows that the beam size remains stable across the full range of insertion device gaps, which is the correct physical outcome. The framework carries out the entire experiment automatically: it finds the right control variables, pulls historical data to define the scan range, generates a safe measurement script, runs the scan, and produces an annotated plot. From a single natural-language request, the system completes a multi-step procedure that normally requires an expert hours to set up.
Figure 3: Hysteresis measurement of insertion device gap versus vertical beam size at the ALS. The execution plan generated by the Alpha Berkeley Framework combined historical range extraction, automated script generation, and real-time machine control. The agent performed a 30-point bidirectional gap sweep with 5 repeated measurements per point, producing the plot shown here for one device.
@@ V Conclusion
We have presented the Alpha Berkeley Framework, a distributed multi-agent architecture designed for complex, information-driven workflows in heterogeneous scientific environments. By dynamically assembling planning agents from capability-specific modules, the framework achieves both modularity and scalability. It supports the generation and sharing of intermediate artifacts, such as code and structured data, that can be consumed by other agents, enabling flexible orchestration and emergent workflows. Human oversight is embedded as a first-class feature, with planning mode and structured interrupts ensuring safety and transparency in high-stakes operations.
The framework has been successfully demonstrated at the ALS, where it executed a multi-step analysis and control procedure directly on the accelerator. This deployment illustrates how agentic systems can enhance the usability, reliability, and efficiency of facility operations by translating natural language instructions into reproducible experimental workflows.
Looking ahead, the framework is being applied beyond accelerator operations to beamlines at the ALS, where it is used to integrate data acquisition, analysis, and control. These ongoing efforts highlight its readiness for routine deployment in production environments and its versatility as a facility-wide capability for scientific user environments.

### v2

@@ IV.1 Wind Farm Example
Consider the following user request:
@@
This example shows how our framework automatically decomposes this query into a coordinated execution plan that transforms this user request into a transparent, interpretable workflow of steps that are using autonomous agents for temporal reasoning, operating with diverse data sources, using domain expertise and even performing data analysis, all while maintaining the transparency and human oversight necessary for reliable industrial applications.
@@
It demonstrates the framework’s ability to coordinate heterogeneous data sources within a unified workflow. In the workflow, the Steps 2-4 retrieve turbine sensor data, weather measurements, and extract performance benchmarks from technical documentation. The knowledge retrieval step (Step 4) highlights the framework’s ability to integrate external data sources, incorporating databases and knowledge repositories via the data source provider pattern. The knowledge provider retrieves numerical parameters and structured data that are formatted for direct consumption by downstream Python code generation, showcasing the integration of external knowledge systems into dynamic analysis workflows.
To reliably generate Python code, the system first performs automated hierarchical analysis planning in response to a user’s query, followed by dynamic code generation. The language model initially produces a structured computational plan that divides the analysis into distinct stages such as data preparation, performance metric calculation, and benchmark comparison. This plan is then translated into executable Python code that correctly accesses and uses the structured context data produced in the preceding steps.
@@
IV.2 Advanced Light Source: Production Validation of the Framework
The Advanced Light Source (ALS) at Lawrence Berkeley National Laboratory is a U.S. Department of Energy synchrotron facility that has provided bright X-ray and ultraviolet light to a broad scientific community for more than three decades. To validate the Osprey Framework under real operational conditions, it was integrated into the ALS control room, where the accelerator is operated and optimized to deliver beams to more than 40 experimental stations.
This deployment, referred to as the Accelerator Assistant, serves as a production-scale test of the framework’s architecture. It demonstrates that the same design principles - dynamic capability classification, plan-first orchestration, and sandboxed execution - remain robust within the strict safety and reliability constraints of a running accelerator. Additional information on the accelerator-specific integration and operational use at the ALS is available in Ref. Hellert et al. (2025).
@@ IV.2 Advanced Light Source: Production Validation of the Framework
The outcome of this automated procedure is summarized in Figure 3, which shows the hysteresis measurement of vertical beam size as a function of insertion device gap. Following a single natural language request, the framework plans and executes the complete sequence autonomously, identifying the relevant control variables, retrieving historical data to define the scan range, generating a safe measurement script, performing the bidirectional sweep, and producing an annotated plot, all without any need for operator steering. Physically, the observed stability of the beam size across the full gap range confirms correct machine behavior. From one request, the system performs a multistage experimental procedure including historical analysis, parameter scan, and visualization that would normally take an expert operator several hours to prepare and complete manually.
Figure 3: Representative hysteresis measurement of insertion device gap versus vertical beam size at the ALS. The execution plan generated by the framework combined historical range extraction, automated script generation, and real-time machine control. The agent performed a 30-point bidirectional gap sweep with five repeated measurements per point, validating the framework’s capability for safe end-to-end orchestration in a live accelerator environment.
@@ V Conclusion
We have presented the Osprey Framework, a domain-agnostic, distributed multi-agent architecture designed for complex, information-driven workflows across diverse safety-critical environments. By dynamically assembling planning agents from capability-specific modules, the framework achieves both modularity and scalability across different domains. It supports the generation and sharing of intermediate artifacts, such as code and structured data, that can be consumed by other agents, enabling flexible orchestration and emergent workflows. Human oversight is embedded as a first-class feature, with planning mode and structured interrupts ensuring safety and transparency in high-stakes operations.
The framework has been successfully deployed in the production environment of the Advanced Light Source, where it autonomously executed a complete analysis and control procedure directly on the accelerator. This deployment demonstrates its practical utility in live facility operations and shows how agentic systems can improve usability, reliability, and efficiency by translating natural language instructions into reproducible experimental workflows.
Looking ahead, the framework’s domain-agnostic architecture positions it for broader adoption across scientific instruments beyond accelerator operations. Current applications extend to beamline operations at the ALS, integrating sample positioning, imaging, and experimental data analysis. These ongoing efforts highlight both its readiness for routine deployment in production environments and its adaptability across diverse safety-critical domains requiring coordinated workflow orchestration.